Please help me solve this application of systems of linear equations problem:
Each course at a community college is worth 1 or 2 credits. A group of students is taking a total of 26 courses that are worth a total of 39 credits. How many 1- credit courses are being taken? How many 2- credit courses are being taken?
I know that I need a system of linear equations. I made x equal the number of 1 credit courses and y equal the number of 2 credits courses. The first linear equation I made from the information provided is x + y= 39, but that's as far as I got.

Found 2 solutions by ankor@dixie-net.com, cahlove:
Answer by ankor@dixie-net.com(22740)

Each course at a community college is worth 1 or 2 credits.
A group of students is taking a total of 26 courses that are worth a total of 39 credits.
How many 1- credit courses are being taken?
How many 2- credit courses are being taken?
:
Your first equation should be:
x + y = 26
The 2nd equation should be:
1x + 2y = 39

Answer by cahlove(1)

Your first equation should be:
x + y = 26
The 2nd equation should be:
1x + 2y = 39
Now that we know the two equations we can use the addition/ elimination method to find our answer.
Let's use the first equation, since the coefficients are 1.
We will multiply the first equation by -2 so the Y variable will cancel out.
-2(x + y = 26)
-2x - 2y = -52
Now that the Y variables are the same, we can add equation two and the new equation we made.
-2x - 2y = -52
+ 1x + 2y = 39
-1x = -13
x = 13
Now that we have x we can plug in back into the original equations to find the value of the y variable. We will use the first equation since it is easier.
x + y = 26
13 + y = 26
y = 13
To check your work, plug x = 13 and y = 13 back into equations 1 and 2.
Equation 1
x + y = 26
13 + 13 = 26
26 = 26
Equation 2
1x +2y = 39
1(13) + 2(13) = 39
13 + 26 = 39
39 = 39
How many 1- credit courses are being taken? Thirteen 1- credit courses were taken
How many 2- credit courses are being taken? Thirteen 2- credit courses were taken